OpenStudy (-welp-):

How do I find the solution set? x-2/x+4 = 2- 4/x

OpenStudy (-welp-):

\[\frac{ x-2 }{ x+4 } = 2 - \frac{ 4 }{ x }\]

OpenStudy (freckles):

first make notice x+4 cannot be zero...we also must not have x be 0 otherwise those fractions bill be zero on bottom anyways try getting rid of the fractions by multiplying both sides by x(x+4)

OpenStudy (-welp-):

I'm not sure what I'm supposed to do after -x^2-10x+16 I know the solution set is {-8,2} (I used mathway)

OpenStudy (freckles):

do you mean -x^2-10x+16=0 how did you get -10x?

OpenStudy (-welp-):

\[ \frac{ x-2 }{ x+4 } * x(x+4) = 2* x(x+4) - \frac{ 4 }{ x } * x(x+4) \] \[x(x-2)= 2x(x+4) - 4(x+4)\] \[x^2 -2x= 2x^2+8x-4x-16\] \[x^2-6x=2x^2-16\] \[x^2-6x+16=2x^2\] \[-x^2-6x+16\] This right?